# zip2zip: Inference-Time Adaptive Vocabularies for Language Models via Token Compression

### v1

@@ 2.5 Efficiency Advantage
zip2zip improves efficiency by increasing the average token length, thereby reducing the number of tokens required to represent the same text. This compression applies to both inputs (e.g., prompts) and outputs (e.g., completions), leading to shorter effective context lengths. As a result, the model performs fewer computations—both in the attention mechanism and the feedforward layers—and, more importantly, requires fewer autoregressive decoding steps during inference. Since the latency of large language models is primarily driven by the cost of sequential decoding, reducing the number of output tokens by $n$leads to an approximate $n$speedup in decoding latency, which we will demonstrate empirically in Section 3.6. A more detailed discussion of FLOPs is provided in Appendix B for completeness.
@@ 3 Experiments
To evaluate the effectiveness of zip2zip, we adapt the Phi-3 models (3B and 14B) within the zip2zip framework. We evaluate our adapted models across four dimensions: (1) token efficiency, (2) language modeling perplexity, (3) downstream task performance, (4) inference efficiency. For perplexity and downstream benchmarks, we use the widely adopted lm-evaluation-harness framework (Gao et al., 2024).
3.1 Training Setup
Rather than updating the full model weights, we adopt parameter-efficient finetuning using LoRA (Hu et al., 2022). In addition, we train the hyper-embedding and hyper-projection modules. We set the maximum merge size to $M=3$ and use a two-layer transformer encoder as the hyper-encoder. Ablation studies on $M$ and hyper-encoder architecture can be found in Appendix A. For comparison, we also perform continual finetuning of the base model using LoRA under identical training conditions, serving as a baseline (denoted as Cont. Finetune in the Tables) The fine-tuning process is highly efficient, requiring approximately 10 H100-GPU hours for a 4B-parameter model and up to 40 H100-GPU hours for a 14B-parameter model, using only 0.1 billion training tokens. Interestingly, the reconstruction loss converges to near zero during training, indicating that the model can almost perfectly recover the original base token sequences from the hypertoken representations. This highlights the learned compression is highly information-preserving. Details of the training setup, compute infrastructure, and dataset curation are provided in Appendices D and E.
3.2 Sample Outputs and Hypertoken Patterns
We present several examples to provide intuition into how the zip2zip model generates text.
Figure 3: Zip2Zip output examples. Blue: base tokens; Yellow: hypertokens (composed of 2 base tokens); Orange: hypertokens (composed of 3+ base tokens).
We see that the model successfully generates a mixture of hypertokens and base tokens in the output (see Figure 3). The hypertoken ratio is as high as 40% in the Python code generation example, and 20% in the biomedical text generation example. Many of the hypertokens correspond to semantically meaningful units or domain-specific terms as shown in Table 1. For a more fine-grained visualization of hypertoken with zip2zip , we provide visualizations of token streams in Appendix 8.
Table 1: Examples of hypertokens formed by zip2zip across three domains
@@ 3.3 Token Efficiency
Given an input text $x$ and a tokenizer, we define token efficiency $\eta:=\frac{\mathrm{Bytes}(x)}{\mathrm{Tokens}(x)}$ as the average number of bytes represented by each token, where $\mathrm{Bytes}(x)$ refers to the number of bytes in the UTF-8 encoding of $x$. This measures how compactly a tokenizer encodes input text—higher values of $\eta$ indicate more efficient tokenization.
We evaluate token efficiency using the tokenizers of four LLMs—Llama-3 (Grattafiori et al., 2024), Qwen-2 (Yang et al., 2024), Phi-4 (Abdin et al., 2024), and Gemma-3 (Team et al., 2025)—each associated with a different base vocabulary size ranging from 128K to 256K. Token efficiency is measured across five representative domains, sampled from publicly available datasets: code (Lozhkov et al., 2024b), math (LI et al., 2024), chat (Ding et al., 2023), multilingual (Penedo et al., 2024), and web (Lozhkov et al., 2024a). Table 2 shows that applying LZW zip2zip consistently improves token efficiency across all tokenizer and domains. Gains are particularly strong in structured domains like code and math—50% higher than the base tokenizer. Interestingly, models with larger vocabulary sizes do not always achieve better token efficiency, suggesting that simply enlarging the vocabulary size is not sufficient to improve it.
Table 2: Token efficiency ($\mathrm{bytes}/\mathrm{token}$) across domains for different tokenizers w/wo zip2zip.
@@ 3.4 Perplexity
Table 3: Byte-perplexity (↓) on four corpora using a 1024-token context window.
Model Method Wiki Pile mC4 dC4
Phi-3.5-4B Base 1.58 1.79 1.88 1.74
Cont. finetune 1.59 1.81 1.88 1.74
zip2zip 1.69 1.95 2.00 1.82
Phi-3-14B Base 1.43 1.72 1.82 1.67
Cont. finetune 1.47 1.79 1.86 1.68
zip2zip 1.56 1.90 1.96 1.75
[…]
@@ 3.5 Evaluation on NLP Benchmarks
Table 4: Two-shot accuracy (in %) across 7 NLP benchmarks, higher is better. Standard deviations (bootstrapped) $\approx$ 0.02 across all tasks.
Model Method ARC-c ARC-e HS OBQA PIQA WG GSM8K
Phi-3.5-4B Base 0.60 0.83 0.66 0.46 0.79 0.75 0.82
Cont. finetune 0.60 0.82 0.63 0.47 0.82 0.75 0.40
zip2zip 0.57 0.83 0.61 0.46 0.82 0.75 0.15
Phi-3-14B Base 0.62 0.80 0.70 0.51 0.83 0.76 0.84
Cont. finetune 0.62 0.88 0.66 0.52 0.87 0.80 0.52
zip2zip 0.62 0.86 0.68 0.51 0.85 0.79 0.25
We next evaluate zip2zip’s performance on real-world tasks. We evaluate on seven widely used NLP benchmarks, including ARC-[Challenge, Easy] (Clark et al., 2018), HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019) and GSM8K (Cobbe et al., 2021). As shown in Table 4, the model finetuned with zip2zip performs similarly to the baseline on most tasks. However, on GSM8K, where the primary task involves numerical computation, the model exhibits significant degradation. Due to the sensitivity of such tasks to tokenization, it occasionally generates malformed or repeated numbers. While token-level operations are already known to be challenging for LLMs (Singh and Strouse, 2024), adaptive tokenization appears to exacerbate this issue.
To validate the effectiveness of zip2zip on non-English languages, we evaluate the model on machine translation tasks, including WMT14 (Macháček and Bojar, 2014), WMT16 (Bojar et al., 2016). The results, shown in Table 5, indicate a small performance degradation across BLEU, CHRF, and TER metrics when using zip2zip. However, the drop is relatively minor, suggesting that the model retains strong multilingual capabilities even in the compressed representation.
Table 5: Machine translation performance on WMT benchmarks. Scores are averaged across both translation directions. Standard deviations (approximately $1.0\sim 2.0$) are reported in Table 10 in Appendix C.
@@ 3.6 Inference Efficiency
zip2zip reduces decoding time by lowering the number of tokens that need to be generated. However, it introduces additional FLOPs due to the on-the-fly computation of hyper-embeddings by the hyper-encoder. To address this overhead, we implement hyper-embedding caching and optimize the computation using a custom Triton kernel. We report separate timings for prefill and decoding across multiple models, with and without zip2zip, in Table 6.
Table 6: Throughput (tokens/sec) comparison of the zip2zip framework against the baseline Hugging Face Transformers generate and MLX generate implementation. Performance is detailed for prefill and decode phases across various context lengths (first value in column headers) combined with a 256-token generation length. zip2zip demonstrates notable throughput improvements, in both prefill and decoding phase.
@@
Figure 4: zip2zip tokenizer latency (ms) vs. HF tokenizer.
As we show in Table 6, zip2zip achieves a significant speedup in all four settings. Both prefill and decoding times are significantly reduced, with the most substantial gains observed in the 512+256 setting with the Phi-3.5-4B model. Improvements are significantly stronger on datacenter-grade GPUs like the NVIDIA H100 and more modest on consumer hardware (e.g., Apple M1).
Efficient LZW Tokenization. zip2zip introduces an additional LZW compression step during inference and a decompression step at the end of generation. As a result, the efficiency of LZW-integrated tokenization is important to overall performance. To minimize overhead, we implemented a Rust-based zip2zip tokenizer that outperforms the Python version (see Figure 4) and matches the latency of HuggingFace’s fast BPE tokenizer.
@@ 4 Related Work
Vocabulary Expansion. Several works have explored expanding the tokenizer vocabulary to better support specific domains or languages. Zhao et al. (2024); Kim et al. (2024); Liu et al. (2023, 2024) adapt LLaMA to Chinese, Korean, and specialized domains such as mental health and law by appending new tokens. Wang et al. (2025); Liu et al. (2024) conducted studies on how to effectively expand the vocabulary by better selecting the subset of tokens to add. In contrast, zip2zip is the first to enable dynamic vocabulary expansion at inference time, constructing new tokens based on the input context without requiring retraining or modifying the tokenizer ahead of time.
Prompt Compression. Prompt compression methods include GistTokens (Mu et al., 2023), Selective Context (Li et al., 2023), LLMLingua (Jiang et al., 2023), Summary Vectors (Chevalier et al., 2023), In-context Autoencoder (Ge et al., 2024), and others (Wingate et al., 2022) reduce the input token length and but do not impact the number of output tokens, which often dominates overall generation time. In contrast, zip2zip compresses both the input and output token sequences.
Latent Tokens Representation. The concept of latent token representations, or patches, has been mostly explored in computer vision, with methods like Token Merging (Bolya et al., 2023) and Token Pooling (Marin et al., 2023) aiming to reduce sequence length while preserving semantic content. Recently, Byte Latent Transformer (BLT) (Pagnoni et al., 2024) extended this concept to language modeling by discarding tokens entirely and operating directly at the byte level. Both BLT and zip2zip adopt a hierarchical modeling of input for LLMs, but they differ in three key ways: (1) Goal: BLT aims to replace the tokenizer, whereas zip2zip seeks to expand and improve it; (2) Algorithm: BLT uses entropy-based segmentation, while zip2zip applies LZW-based token compression; (3) Training: BLT requires training from scratch, whereas zip2zip enables continued adaptation of pretrained models. Lester et al. (2024) propose improving language model efficiency by training LLMs directly on text compressed with arithmetic coding. While both approaches leverage compression to enhance efficiency, zip2zip emphasizes dynamic vocabulary expansion to enable uptraining of existing models. In contrast, Lester et al. (2024) requires training from scratch.
@@ 5 Discussion and Limitations
Codebook Management Strategy. The LZW algorithm grows the codebook linearly with the number of tokens in the context window. Empirical results show that only about 25% of hypertokens are reused during generation, leaving substantial room for optimization. Two potential improvements are: (1) pruning or selective retention strategies to reduce unused entries, and (2) codebook prefilling, which could be beneficial if likely tokens can be anticipated before input processing.
Compression–Quality Trade-off. There is an inherent trade-off between compression and modeling: as the token space is compressed more aggressively, redundancy is reduced—but so is predictability—making it harder for the model to forecast the next (hyper)token. In the extreme, optimal compression schemes such as arithmetic coding produce sequences that are statistically indistinguishable from random noise, rendering them unlearnable by language models (Lester et al., 2024). Empirically, we observe this effect as increased perplexity under higher compression levels (Table 3), which can undermine the benefits of compression by degrading generation quality (though minor in the tasks in Table 4 and Table 5). Striking the right balance between compression and model performance remains an important direction for future research.

### v2

@@ 2.5 Efficiency Advantage
zip2zip improves efficiency by increasing the average token length, thereby reducing the number of tokens required to represent the same text. This compression applies to both inputs (e.g., prompts) and outputs (e.g., completions). As a result, the model performs fewer computations—both in the attention mechanism and the feedforward layers—and, more importantly, requires fewer autoregressive decoding steps during inference. Since the latency of large language models is primarily driven by the cost of sequential decoding, reducing the number of output tokens by nn% leads to an approximate nn% speedup in decoding latency, which we will demonstrate empirically in Section 3.6. A more detailed discussion of FLOPs is provided in Appendix E for completeness.
2.6 Entropy Invariance under Lossless Compression
Before turning to empirical results, we analyze whether a lossless compression of the data representation can fundamentally alter the achievable performance of a model. We show that for any lossless mapping $g$, there always exists a corresponding transported model distribution in the compressed space that achieves exactly the same (cross-)entropy as in the original space.
Let $\mathcal{X}$ be the original alphabet and $\mathcal{Z}$ an arbitrary alphabet obtained via a lossless compressor $g:\mathcal{X}^{\ast}\to\mathcal{Z}^{\ast},$ which is a bijection onto its image. Denote by $P_{X}$ the true distribution over sequences $x\in\mathcal{X}^{\ast}$, and by $p_{\theta}$ the model distribution on the same space. The corresponding push-forward (compressed-space) true distribution $P_{Z}$ and model distribution $p_{\gamma}$ are defined as
$P_{Z}(z)=P_{X}(g^{-1}(z)),\qquad p_{\gamma}(z)=p_{\theta}(g^{-1}(z)),\qquad z\in\mathcal{Z}^{\ast}.$
Theorem 2.1 (Entropy invariance under lossless compression).
If $g$ is lossless (i.e., bijective onto its image), then the total entropy and cross-entropy are invariant under the transformation:
$H(P_{Z})=H(P_{X}),\qquad H(P_{Z},p_{\gamma})=H(P_{X},p_{\theta}).$
A detailed proof is provided in Appendix G.
The theorem implies that the optimal achievable cross-entropy in the compressed representation is identical to that in the original domain: for any model family on $\mathcal{X}^{\ast}$, one can always construct a corresponding model family on $\mathcal{Z}^{\ast}$ via push-forward that attains the same likelihood. In practice, the training process can be viewed as an attempt to approximate this transported model through optimization; however, convergence to the target model is not guaranteed (see Section 5).
@@ 3 Experiments
To evaluate the effectiveness of zip2zip, we perform continued pretraining on the Phi-3 models (3B and 14B) within the zip2zip framework. We train a single model on a general-purpose corpus and evaluate it across four dimensions: (1) token efficiency, (2) language modeling perplexity, (3) downstream task performance, and (4) inference efficiency. This setup allows us to assess how well zip2zip generalizes to diverse domains without any task- or domain-specific fine-tuning. For perplexity and downstream benchmarks, we use the widely adopted lm-evaluation-harness framework (Gao et al., 2024).
Figure 4: Phi-3.5-zip2zip output examples. Blue: base tokens. Yellow: hypertokens (composed of 2 base tokens). Orange: hypertokens (composed of 3+ base tokens).
Table 1: Examples of hypertokens formed by Phi-3.5-zip2zip across three domains
@@
3.1 Training Setup
Rather than updating the full model weights, we adopt parameter-efficient finetuning using LoRA (Hu et al., 2022). In addition, we train the hyper-embedding and hyper-unembedding modules. We set the maximum merge size to $M=3$ and use a two-layer transformer encoder as the hyper-encoder. The loss weighting coefficient $\lambda$ was chosen to be $0.1$, as justified in Appendix I. Ablation studies on $M$ and the hyper-encoder architecture can be found in Appendix C. For comparison, we also perform continual pretraining of the base model using LoRA under identical training conditions, serving as a baseline (denoted as Cont. Pretraining in the tables). The continual pretraining process is highly efficient, requiring approximately 10 H100-GPU hours for a 4B-parameter model and up to 40 H100-GPU hours for a 14B-parameter model, using only 0.1 billion training tokens. Interestingly, the reconstruction loss converges to near zero during continued pretraining, indicating that the model can almost perfectly recover the original base token sequences from the hypertoken representations. This highlights the learned compression is highly information-preserving. Details of the training setup, compute infrastructure, and dataset curation are provided in Appendices I and J.
3.2 Qualitative Examples and Hypertoken Patterns
We present several examples (Figure 4 and Table 1) to provide intuition into how the zip2zip model generates text. We see that the model generates a mixture of hypertokens and base tokens in the output (Figure 4). The hypertoken ratio is as high as 40% in the Python code generation example, and 20% in the biomedical text generation example. Many of the hypertokens correspond to semantically meaningful units or domain-specific terms as shown in Table 1. For a more fine-grained visualization of hypertoken with zip2zip, we provide visualizations of token streams in Figure 10 in the appendix.
@@ 3.3 Token Efficiency
Given an input text $x$ and a tokenizer, we define the token efficiency $\eta:=\frac{\mathrm{Bytes}(x)}{\mathrm{Tokens}(x)}$ as the average number of bytes represented by each token (also called compression ratio), where $\mathrm{Bytes}(x)$ refers to the number of bytes in the UTF-8 encoding of $x$. This measures how compactly a tokenizer encodes input text—higher values of $\eta$ indicate more efficient tokenization. We evaluate token efficiency using the tokenizers of four LLMs—Llama-3 (Grattafiori et al., 2024), Qwen-2 (Yang et al., 2024), Phi-4 (Abdin et al., 2024), and Gemma-3 (Team, 2025)—each associated with a different base vocabulary size ranging from 128K to 256K. Token efficiency is measured across five representative domains, sampled from publicly available datasets: code (Lozhkov et al., 2024b), math (LI et al., 2024), chat (Ding et al., 2023), multilingual (Penedo et al., 2024), and web (Lozhkov et al., 2024a). Table 2 shows that applying LZW zip2zip consistently improves token efficiency across all tokenizer and domains. Gains are particularly strong in structured domains like code and math—with gains of 48% and more over the base tokenizer. Interestingly, models with larger vocabulary sizes do not always achieve better token efficiency, suggesting that simply enlarging the vocabulary size is not sufficient to improve it.
Table 2: Token efficiency (bytes per token) across domains for different tokenizers with and without zip2zip.
@@ 3.4 Perplexity
We evaluate the perplexity of zip2zip models on four corpora: Wikitext (Merity et al., 2016), The Pile (Gao et al., 2020), and two subsets of Paloma (Magnusson et al., 2023): mC4, a multilingual subset of C4, and dC4 (aka C4-100D), a subset of C4 spanning 100 domains. Given a token sequence $x=x_{1},\dots,x_{N}$, and a model $q$, perplexity and byte-level perplexity (Radford et al., 2019; Magnusson et al., 2023) are defined as: $\text{PPL}:=\left(\prod_{i=1}^{N}q(x_{i})\right)^{-1/N}$, $\text{Byte-PPL}:=\left(\prod_{i=1}^{N}q(x_{i})\right)^{-1/B}={\text{PPL}}^{1/\eta}$, where $B$ is the number of UTF-8 bytes of the text, and $\eta$ denotes the token efficiency (i.e., bytes per token). Token-level perplexity depends on the tokenization scheme and is unsuitable for cross-tokenizer comparison. We instead report byte-level perplexity, a vocabulary-agnostic metric that normalizes for tokenization differences. Table 3 (right panel) shows that zip2zip models see a modest increase in byte-level perplexity, indicating a slight drop in language modeling performance.
Table 3: Two-shot accuracy across seven NLP benchmarks (left) and byte-level perplexity (↓) on four corpora using a 1024-token context window (right). Standard deviations (bootstrapped) $\approx$ 0.02 across all tasks.
Model Method ARC-c ARC-e HS OBQA PIQA WG GSM8K Wiki Pile mC4 dC4
Phi-3.5-4B Base 0.60 0.83 0.66 0.46 0.79 0.75 0.82 1.58 1.79 1.88 1.74
Cont. pretrain 0.60 0.82 0.63 0.47 0.82 0.75 0.40 1.59 1.81 1.88 1.74
zip2zip 0.57 0.83 0.61 0.46 0.82 0.75 0.15 1.69 1.95 2.00 1.82
Phi-3-14B Base 0.62 0.80 0.70 0.51 0.83 0.76 0.84 1.43 1.72 1.82 1.67
Cont. pretrain 0.62 0.88 0.66 0.52 0.87 0.80 0.52 1.47 1.79 1.86 1.68
zip2zip 0.62 0.86 0.68 0.51 0.85 0.79 0.25 1.56 1.90 1.96 1.75
@@ 3.5 Evaluation on NLP Benchmarks
We next evaluate zip2zip’s few-shot performance on real-world tasks. We evaluate on seven widely used NLP benchmarks, including ARC-[Challenge, Easy] (Clark et al., 2018), HellaSwag (Zellers et al., 2019), LAMBADA (Paperno et al., 2016), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019) and GSM8K (Cobbe et al., 2021). As shown in Table 3, the model continued-pretrained with zip2zip performs similarly to the baseline on most tasks. However, on GSM8K, where the primary task involves numerical computation, the model exhibits significant degradation. While token-level operations are already known to be challenging for LLMs (Singh and Strouse, 2024), it is possible that adaptive tokenization exacerbates this effect, though further validation is required to confirm this hypothesis.
Multilinguality.
To validate the effectiveness of zip2zip on non-English languages, we evaluate the model on machine translation tasks, including WMT14 (Macháček and Bojar, 2014), WMT16 (Bojar et al., 2016). The results, shown in Table 4, indicate a small performance degradation across the BLEU, CHRF, and TER metrics when using zip2zip. However, the drop is relatively minor, suggesting that the model retains strong multilingual capabilities even in the compressed representation. Additional experiments on multilingual QA benchmarks are provided in Appendix H.2.
Table 4: Machine translation performance on WMT benchmarks. Scores are averaged across both translation directions. Standard deviations (approximately $1.0\sim 2.0$) are reported in Table 11 in Appendix H.
@@ 3.6 Inference Efficiency
zip2zip reduces decoding time by lowering the number of tokens that need to be generated. However, it introduces additional FLOPs due to the on-the-fly computation of hyper-embeddings by the hyper-encoder. To address this overhead, we implement hyper-embedding caching and optimize the computation using a custom Triton kernel. We report separate timings for prefilling and decoding across multiple models, with and without zip2zip, in Table 5. As we show in Table 5, zip2zip achieves a significant speedup in all four settings. Both prefilling and decoding times are significantly reduced, with the most substantial gains observed in the 512+256 setting with the Phi-3.5-4B model. Improvements are significantly stronger on datacenter-grade GPUs like the NVIDIA H100 and more modest on consumer hardware (e.g., Apple M1).
Table 5: Throughput (tokens/sec) comparison of the zip2zip framework against the baseline HuggingFace Transformers generate and MLX generate implementation. Performance is detailed for prefilling and decode phases across various context lengths (first value in column headers) combined with a 256-token generation length. zip2zip demonstrates notable throughput improvements, in both the prefilling and decoding phases.
@@
Efficient LZW-Tokenizer Implementation. zip2zip introduces an additional LZW compression step during inference and a decompression step at the end of generation. As a result, the efficiency of LZW-integrated tokenization is important to overall performance. To minimize overhead, we implemented a Rust-based zip2zip tokenizer that outperforms the Python version (see Figure 7) and matches the latency of HuggingFace’s fast BPE tokenizer.
@@ 4 Related Work
Domain-Adapted Tokenizers. Several works have explored tokenizer adaptation by expanding the token vocabulary to better support specific domains or languages. Zhao et al. (2024); Kim et al. (2024); Liu et al. (2023, 2024a) adapt the Llama tokenizer to Chinese, Korean, and specialized domains such as mental health and law by adding new tokens. However, these approaches yield a fixed vocabulary that does not adapt during inference.
Input Compression for LLMs. Prompt compression methods such as gist tokens (Mu et al., 2023), selective context (Li et al., 2023), LLMLingua (Jiang et al., 2023), summary vectors (Chevalier et al., 2023), in-context autoencoders (Ge et al., 2024), and others (Wingate et al., 2022) reduce the context length by lossy compression. While their lossy compression nature enables high compression ratios, these prompt compression methods can only compress input tokens, but not the output tokens, although output tokens typically dominate generation time under low-batch workloads. Lester et al. (2024) propose improving language model efficiency by training LLMs directly on text compressed with arithmetic coding.
Transformers with Dynamic Embeddings. Architecture-wise, zip2zip employs a dynamic embedding layer built upon transformer blocks. Similar ideas have been explored in prior work aimed at reducing the computational cost of transformers, including Hourglass (Nawrot et al., 2022), dynamic-pooling transformer (Nawrot et al., 2023), MegaByte (Yu et al., 2023), Toucan (Fleshman and Durme, 2023), Learn-Your-Token (Thawani et al., 2023), SpaceByte (Slagle, 2024), ZeTT (Minixhofer et al., 2024), dynamic tokenization (Feher et al., 2025), BLT (Pagnoni et al., 2025), and H-Net (Hwang et al., 2025). These approaches vary in their model architectures and chunking strategies. Dynamic Vocab (Liu et al., 2024b) is probably the closest in terms of conceptual motivation, as it also expands the vocabulary dynamically during generation. The main difference lies in the dynamic vocabulary construction algorithm and the model training procedure.
@@ 5 Discussion and Limitations
Codebook Management Strategy. The LZW algorithm grows the codebook linearly with the number of tokens in the context window. Empirical results show that only about 25% of hypertokens are reused during generation, leaving substantial room for optimization. Two potential improvements are (1) pruning or selective retention strategies to reduce unused entries, and (2) codebook prefilling, which could be beneficial if likely tokens can be speculated ahead of input processing.
Optimization Under Lossless Compression. Since zip2zip employs lossless compression, the achievable performance is theoretically invariant under the transformation: a transported model, as described in Section 2.6, can attain identical likelihood to that in the original space. Empirically, however, we observe a mild increase in perplexity under compression (Table 3), indicating that the trained model does not perfectly recover the transported model. This discrepancy arises from optimization challenges rather than representational limits—gradient descent may converge more slowly or settle in suboptimal regions due to more complex loss landscape. Understanding this optimization difficulty—how the search landscape changes under compression and whether targeted preconditioning or extended training budgets can close the gap—remains an important question for future work.